On a measure with a mean of 10 and a standard deviation of 2, Sumathi scored 13. What proportion of people score farther from the mean than Sumathi? (Give your answer to at least 3 places past the decimal point) The answer is not 0.0668
Answer #1

Solution :

P(x > 13) = 1 - P(x < 13)

= 1 - P[(x - ) / < (13 - 10) / 2]

= 1 - P(z < 1.5)

= 1 - 0.9332

= 0.067